The salaries of the employees in a company follow the normal distribution with mean $16000 and standard deviation $800.
(a) What is the probability that the salary of an employee is higher than $18000?
(b) There is 80% chance that the salary of an employee is less than $t. Find the value of t.
(c) If 20 employees are randomly chosen from the company and the mean salary of the employees is calculated. The manager of the company claims that over 20% of the mean salaries of the employee are higher than $16200. Do you agree?

Answer by Boreal(15235)

z=(x-mean)/sd
z > (18000-16000)/800 or 2.5
that probability is 0.0062
-
z (0.80)=0.8416
0.8416=(t-16000)/800
673.29=t-16000, will round to nearest integer
t=$16,673.
-
z=(16200-16000)/800/sqrt(20)
-200*sqrt(20)/800=.
=1.118
That probability of z> than that value is 13.2%
The sd of the sampling distribution with n=20 has mean $16000 as before, but sd is 178.88